AIDA GELINA BRIKEN nToF CRIB ISOLDE CIRCE nTOFCapture DESPEC DTAS EDI_PSA 179Ta CARME StellarModelling DCF K40
  CARME, Page 18 of 33  ELOG logo
ID Date Author Subject Text Attachments
  319   Mon Feb 28 02:16:00 2022 JM, OFCurrent status03:21

Temperature setpoint 110 K

Rutherford rate stable. Min peak
 6x 
  318   Mon Feb 28 01:16:02 2022 JM, OFTemperature decrease 115K to 110K02:15

Density remains stable at
155 K
 20220228_021455.jpg20220228_023832.jpg20220228_025416.jpg 
  317   Mon Feb 28 00:24:05 2022 JM, OFCurrent status01:23

Temperature setpoint = 115 k

Rutherford rates still appears
 7x 
  316   Sun Feb 27 23:17:33 2022 JM, OFCurrent status00:20

Temperature setpoint 115 K

Rutherford rate appears fairly
 10x 
  315   Sun Feb 27 22:03:51 2022 JM, OFCurrent Status23:00

Target Temperature setpoint 95
K
 9x 
  314   Sun Feb 27 20:38:06 2022 Rui-Jiu Chen, YuriCurrent statusRing current ~5.0e7 particles/s

SW3

- Merger OK
 13x 
  313   Sun Feb 27 20:19:06 2022 Rui-Jiu Chen, Yuriopen the valves of target.    
  312   Sun Feb 27 19:01:02 2022 Rui-Jiu Chen, YuriCurrent statusRing current ~5.96e7 particles/s

SW3

- Merger OK
 13x 
  311   Sun Feb 27 18:46:42 2022 Rui-Jiu Chen, YuriChange target temperaturetime             
Temperature            
E1 current(A)                
  
  310   Sun Feb 27 17:00:45 2022 Rui-Jiu Chen, Yuricurrent statusRing current ~7.76e7 particles/s

SW3

- Merger OK
 13x 
  309   Sun Feb 27 16:30:35 2022 Rui-Jiu Chen, YuriChange target temperature again1. change T from 95 k to 140 k.

2. keep T at 140k for 15 minutes.

3. change T from 140 k to 95 k.
  
  308   Sun Feb 27 15:41:33 2022 Rui-Jiu Chen, Yuricurrent status16:38 change target temperature from 95
k to 140 k.

16:48 change target temperature
 13x 
  307   Sun Feb 27 15:02:16 2022 Rui-Jiu Chen, YuriChange target temperature16:00 change temperature from 95 k to 120k.

16:07 change temperature from 120
k to 95k.
 14x 
  306   Sun Feb 27 14:05:30 2022 Rui-Jiu, RaganCurrent statusRing current ~7.61e7 particles/s

SW3

- Merger OK
 14x 
  305   Sun Feb 27 13:34:24 2022 RSS TDCurrent status14:33: 95K-->140K

14:44: 140K --> 95K

Ring current ~6e7 particles/s
  
  304   Sun Feb 27 12:49:44 2022 TD RSSSunday 27 February 07:00-15:00
13.50 Cryostat set point 120K
 6x 
  303   Sun Feb 27 12:00:13 2022 RSS TDCurrent statusRing current ~7.7e7 particles/s

Merger, Tape server (WS3):
1414 kB/s writing ok
 13-xy-2022-02-27_13-01-36.png13-yrate-2022-02-27_13-01-06.png13-xrate-Screenshot_from_2022-02-27_13-00-50.png 
  302   Sun Feb 27 11:00:08 2022 RSS TDCurrent status12:00: 95K-->120K

12:07: 120K --> 95K

Ring current ~6e7 particles/s
 12-xy-2022-02-27_12-16-19.png12-yrate-2022-02-27_12-15-42.png12-xrate-2022-02-27_12-15-03.png 
  301   Sun Feb 27 10:00:07 2022 RSS TDCurrent statusRing current ~9e7 particles/s

Merger, Tape server (WS3):
1414 kB/s writing ok
 11-xy-2022-02-27_11-01-22.png11-ey-rate2022-02-27_11-00-46.png11-ex-rate-2022-02-27_11-00-27.png 
  300   Sun Feb 27 09:00:09 2022 RSS TDCurrent status10:00: 95K-->120K

10:08: 120K --> 95K

Ring current ~7e7 particles/s
 7x 
ELOG V3.1.4-unknown